If the demand and supply of Chinese coal are
Qd = 60 − 0.5P.
Qs = −2 + 2P.

Determine the new (1) quantity (2) inverse demand price and (3) inverse demand quantity if a
subsidy of $10 is paid on Chinese coal?
1
Expert's answer
2020-11-16T07:34:05-0500
"Q_d=60-0.5(p-10)"


"Q_s=-2+2p"


"60-0.5p+5=-2+2p"


"p=22.8"


"Q=43.6"


"pQ=994.08"